# VLCE: A Knowledge-Enhanced Framework for Image Description in Disaster Assessment

### v1

@@ 6.4 Key Findings and Implications
Exp. Configuration Custom Model LLaVA QwenVL Best Model
ViT_UAV_RescueNet Dataset
1 With Knowledge Graph 272 185 201 Custom
2 Without Knowledge Graph 272 178 195 Custom
ResNet-EuroSAT (xBD Dataset)
3 With Knowledge Graph 640 445 463 Custom
4 Without Knowledge Graph 640 431 448 Custom
Performance Summary
Average Improvement over LLaVA +45.2%
Average Improvement over QwenVL +38.7%
Experiments Won 4/4 0/4 0/4
@@ 7 Case Studies: Impact of Knowledge Graph on Caption Quality
Experimental assessment comprises various models (ViT and QwenVL) configurations run on the RescueNet dataset of 899 test and XBD dataset. We have shown Fig. 19 & Fig.  20 for LLaVA image captioning with KG and without KG for the ViT model based on the RescueNet dataset, tested caption. The figure 19 is provided below on “Remote-Sensing-UAV-image-classification” model, RescueNet Dataset with KG LLaVA Caption.
@@
Therefore, another figure is provided below 20 on “Remote-Sensing-UAV-image-classification” model, RescueNet dataset with out KG LLava caption. This example contains misinformation, repeated words, and incorrect sentences because KG was not used.
@@
We also evaluated the XBD dataset and tested on ResNetEurosat with KG and without KG image captioning. Below is provided a Fig. 21 on ResNetEurosat XBD Dataset with KG LLaVA Caption.
@@
However, our tested fig shows on 22 the ResNeteurosat XBD dataset without KG Lava caption which contains misinformation.
@@
For more evaluation, we also tested on “Remote-Sensing-UAV-image-classification” model, RescueNet dataset and the QwenVL model with KG and without KG analysis. The Fig. 23 is shown below on “Remote-Sensing-UAV-image-classification” RescueNet Dataset with KG QwenVL Caption.
@@
Alternatively, Fig. 24 analysis without KG generates “Remote-Sensing-UAV-image-classification” model, RescueNet Dataset without KG QwenVL Caption with low quality caption.
@@
Moreover, we also tested XBD dataset QwenVL caption with ResNetEurosat output using with KG and without KG caption as an evaluation. The Fig.25 shows XBD Dataset ResNet-EuroSAT Output Using QwenVL with Knowledge Graph.
@@
Furthermore, another evaluation is based on without KG. Below, the Fig.26 shows the caption of XBD Dataset ResNet-EuroSAT Output Using QwenVL without Knowledge Graph with misleading context.

### v3

@@ 6.3.1 CLIPScore Evaluation: Semantic Alignment
Exp. Configuration Custom Model LLaVA QwenVL Best Model
ViT_UAV_RescueNet Dataset
1 With Knowledge Graph 272 185 201 Custom
2 Without Knowledge Graph 272 178 195 Custom
ResNet-EuroSAT (xBD Dataset)
3 With Knowledge Graph 640 445 463 Custom
4 Without Knowledge Graph 640 431 448 Custom
Performance Summary
Average Improvement over LLaVA +45.2%
Average Improvement over QwenVL +38.7%
Experiments Won 4/4 0/4 0/4
@@ 7 Case Studies: Impact of Knowledge Graph on Caption Quality
We experimented with several model configurations and datasets to assess the efficacy of Knowledge Graph (KG) integration. Our investigation evaluates caption quality between KG-enhanced and baseline models utilizing two key datasets: RescueNet and XBD. Several model designs, such as ViT, ResNet-EuroSAT, and other caption generation models (LLaVA and QwenVL), are included in the tests. This section illustrates the distinction between figure captions based on the utilization of KG.
@@
The KG-enhanced caption delivers coherent and contextually precise descriptions. It accurately identifies flooding as the principal type of damage, observes the existence of submerged debris, and deduces recovery operations from the visible vehicles. The narrative maintains a consistent flow without redundancy or fictitious elements.
@@
Conversely, Figure 20 illustrates the baseline model output devoid of KG enhancement, indicating substantial quality deterioration.
@@
The baseline caption demonstrates various failure modes: word duplication (“the the”), nonsensical terms (“stone”) presented without context, (“widespreadvegetation” as a single word), and self-repetition with entire phrases repeating verbatim. These errors illustrate how the lack of organized knowledge results in both syntactic and semantic failures.
We expanded our assessment to the XBD dataset utilizing the ResNet-EuroSAT architecture. Figure 21 illustrates the results enhanced by knowledge graphs.
@@
The KG-enhanced model effectively detects mixed damage scenarios, identifying both forested regions and buildings, while critically differentiating between damaged and undamaged structures. This detailed evaluation is crucial for disaster response planning.
The baseline model devoid of knowledge graph (Figure 22) produces significantly divergent—and concerning—results.
@@
The baseline caption includes fabricated elements (individuals in the field, animal fatalities) and semantically incoherent phrases (“animal deaths may have been temporarily closed”). These fabrications exemplify a significant failure mode in which the model produces credible yet entirely inaccurate information—a particularly perilous consequence for disaster assessment applications.
To evaluate the applicability of KG benefits across various caption generation architectures, we examined the QwenVL model. Figure 23 illustrates the performance of KG-enhanced QwenVL on RescueNet.
@@
The KG-enhanced QwenVL caption exhibits advanced environmental comprehension, accurately recognizing vegetation damage patterns, evaluating ecosystem effects, and deriving pertinent conclusions regarding emergency response implications. The description upholds factual accuracy while offering actionable insights.
Figure 24 demonstrates that the baseline QwenVL output displays analogous degradation patterns to those identified in LLaVA.
@@
The baseline output exhibits significant semantic inaccuracies, including the illogical phrase "food area and causing harm to five people" and fragmented sentences. The fabrication of precise casualty figures is especially alarming, as such misinformation could lead to severe repercussions in genuine disaster response situations. The concluding assessment analyzes QwenVL’s performance on the XBD dataset. Figure 25 illustrates the capabilities enhanced by knowledge graphs.
@@
The KG-enhanced model generates a succinct, precise evaluation that clearly delineates the mixed damage state—essential information for prioritizing rescue and recovery efforts.
In the absence of KG integration (Figure 26), the model’s output is verbose, redundant, and less dependable.
@@
The baseline caption demonstrates phrase redundancy (“showing signs of destruction" is repeated thrice), incomplete phrases (“collapsed trees broken"), and formatting inaccuracies (“areasbuildings" concatenated without spacing). These problems diminish caption effectiveness and may obstruct automated processing systems.
@@
7.1 Summary of Findings
In all assessed configurations, KG integration uniformly enhances caption quality via three principal mechanisms:
  1. 1.
    Elimination of hallucinations: KG-enhanced models prevent the fabrication of details (casualty figures, fictitious individuals) that could mislead disaster response initiatives.
  2. 2.
    Improved coherence: Organized knowledge mitigates redundancy, fragmented sentences, and illogical word amalgamation that afflict baseline models.
  3. 3.
    Enhanced semantic accuracy: KG-grounded captions accurately delineate damage types, disaster categories, and structural conditions, offering actionable intelligence for emergency responders.
These enhancements are consistent across various model architectures (ViT, ResNet-EuroSAT), caption generators (LLaVA, QwenVL), and disaster datasets (RescueNet, XBD), illustrating the generalizability of KG-enhanced methodologies for disaster image captioning.